# TreeEval: Benchmark-Free Evaluation of Large Language Models through Tree Planning

### v1

@@ Abstract
Recently, numerous new benchmarks have been established to evaluate the performance of large language models (LLMs) via either computing a holistic score or employing another LLM as a judge. However, these approaches suffer from data leakage due to the open access of the benchmark and inflexible evaluation process. To address this issue, we introduce TreeEval, a benchmark-free evaluation method for LLMs that let a high-performance LLM host an irreproducible evaluation session and essentially avoids the data leakage. Moreover, this LLM performs as an examiner to raise up a series of questions under a topic with a tree planing strategy, which considers the current evaluation status to decide the next question generation and ensures the completeness and efficiency of the evaluation process. We evaluate $6$ models of different parameter sizes, including $7$B, $13$B, and $33$B, and ultimately achieved the highest correlation coefficient with AlpacaEval2.0 using only around $45$ questions. We also conduct more analysis to show the robustness and reliability of TreeEval. Our code can be accessed via the provided URL [^1] https://github.com/Ashura5/TreeEval.
TreeEval: Benchmark-Free Evaluation of Large Language Models through Tree Planning
Xiang Li East China Normal University xiang.li@stu.ecnu.edu.cn                        Yunshi Lan East China Normal University yslan@dase.ecnu.edu.cn                        Chao Yang Shanghai AI Laboratory yangchao@pjlab.org.cn
1 Introduction
@@
The recent surge in Large Language Models (LLM) has been significant, transitioning from closed-source OpenAI (2023); Team (2023a) to open-source Touvron et al. (2023); et al. (2023a); Jiang et al. (2023) models. Different Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) techniques are further proposed to improve the performance of LLMs Taori et al. (2023); Chiang et al. (2023); Bai et al. (2022); Ouyang et al. (2022); Tunstall et al. (2023a). These LLMs pose the capabilities to solve diverse tasks and are widely used in both academic and industrial fields. Human evaluation is intuitive to assess the performance of LLMs but it is time-consuming and has the risk of including unexpected bias Zheng et al. (2023b); Wang et al. (2024). It becomes vital to investigate the automatic evaluation approaches to evaluate LLMs.
Till now, there are a number of automatic evaluation methods that have been proposed. A line of studies take efforts to annotate some benchmark datasets, such as MMLU, BBH, SocKET, AGIEval, IFEval and HalluQA  Hendrycks et al. (2021); Suzgun et al. (2022); Choi et al. (2023); Zhong et al. (2023); Zhou et al. (2023a); Cheng et al. (2023), to test the different capabilities of an LLM. The evaluated LLM is judged via checking the overlap between the annotated answers and generated answers. In this way, a holistic score is produced to indicate the performance of the LLM. We denote this category of evaluation methods are in a benchmark paradigm. However, the holistic score is inflexible to measure the quality of the outputs from LLMs because the token mismatch unnecessarily indicates the incorrect answer. With the emergence of high-performance LLMs, another line of studies start to leverage them to simulate human evaluation by providing the evaluated LLM with pre-defined benchmark questions and requesting another LLM like GPT-4 to judge its response Zheng et al. (2023a); Li et al. (2023b); Bai et al. (2023); Wang et al. (2023a); Zhang et al. (2023b); Wang et al. (2023c); Li et al. (2023a); Zhu et al. (2023). We denote this category of evaluation methods as LLM-as-judge paradigm. Even though both above lines of methods enable the automatic evaluation with standard pipelines, they encounter severe issue of data leakage. Due to the vast amount of training data used in LLM training, considered a valuable asset by many closed and even open-source models, it is easy to cause benchmark data leakage problems and significantly biases evaluation results.
@@
To solve this issue, we propose a novel evaluation paradigm, which takes an LLM as an examiner to raise questions. The examiner should produce different evaluation session for each time which makes it hard to duplicate the evaluation questions and protect the evaluation benchmark from disclosure for fine-tuning and pre-training an LLM deliberately. However, simply adopting an LLM as examiner would lead to arbitrary evaluation question generation without a goal. Designing such a benchmark-free evaluation method need take the following aspects into consideration: (1) Similar as the question in a benchmark Taori et al. (2023); Zheng et al. (2023a), the generated questions should be derived from certain topics, which ensures the scope of the evaluation. (2) Drawing inspiration from the interview, within a topic, the examiner should generate a line of questions that are diverse to cover different knowledge rather than producing a single question. (3) The generation procedure should be flexible enough to generate mutually connected questions and control the difficulty level of these questions. When the current line of question cannot distinguish two LLMs, more difficult questions should be raised up. Otherwise, the evaluation could be terminated immediately.
To this end, we propose TreeEval, which is a benchmark-free evaluation of LLMs through tree planning. The line of questions within a topic for evaluation are organized in a tree, where each node contains a question. In the process of constructing a tree, we repeatedly revisit the status of the current tree and generate the next node until the tree is enough to differentiate two LLMs. The difference between our evaluation method and previous paradigms can be found in Figure 1. To verify the effect of our method, we evaluate multiple LLMs. The results demonstrate that our method shows similar ranking as AlpacaEval2.0 in LLM-as-judge paradigm with only $45$ questions in average for each round of evaluation. Further analysis shows our advantages in measuring fine-grained capabilities and conducting robust comparison for LLMs.
@@
  • We compare with a set of automatic evaluation baselins, and find that our TreeEval achieves the highest correlation coefficient with AlpacaEval2.0.
2 Related Work
2.1 Methods of LLM Evaluation
Due to the explosive growth and rapid update of LLMs, a significant challenge is to conduct accurate and comprehensive evaluation for them Chang et al. (2023). Early studies leverage open-ended question answering datasets and math word problems as the evaluation benchmarks Touvron et al. (2023); et al. (2023b); Chen et al. (2024) to evaluate the commonsense knowledge and reasoning capabilities of LLMs. Subsequently, more benchmark datasets like MMLU Hendrycks et al. (2021), AGIEval Zhong et al. (2023), IFEval Zhou et al. (2023a) have been elaborately designed to gauge diverse abilities of LLMs. Some studies Wang et al. (2023a, a); Saha et al. (2023) go beyond standard evaluation metrics. They evaluate the quality and accuracy of predicted results through human annotation, which is able to provide a more comprehensive feedback. With the emergence of high-performance LLMs like GPT-4 OpenAI (2023), Gemini Pro Team (2023a), more recent studies start to utilize them to simulate the human evaluation process. In this realm, PandaLM Wang et al. (2023c) strives to provide reproducible and automated comparison between various LLMs by training a LLM as the judge. GPTScore Fu et al. (2023) and G-Eval Liu et al. (2023) utilize GPT-3 and GPT-4 as the judge to evaluate the LLMs with incorporation of in-context learning and chain-of-thought strategies. The above methods rely heavily on a well-organized benchmark dataset while our method is benchmark-free and has LLMs performing as the examiner.
2.2 Data Leakage of LLM Evaluation
@@
3 Methodology
Figure 2: TreeEval system with an illustrative tree for evaluation. The left section contains the components and their workflow in TreeEval. The right section displays a constructed tree within topic Technology and Communication for evaluation (the leaf nodes are shown in red boxes), where each node denotes a question annotated with its topic and evaluation score. We further display the generated questions of the tree in the Appendix 8.6.
3.1 Overall Architecture
Figure 2 depicts the overall architecture of TreeEval. To organize the evaluation in a tree structure, TreeEval incorporates several components including an Examiner, a Judge, and an Eval Controller. After the tree has been constructed, an Aggregator is utilized to aggregate the scores in the tree. The entire framework achieves benchmark-free evaluation of LLMs with a process of tree planning.
In detail, for each session of evaluation, we choose a pair of LLMs for evaluation. The evaluation process starts with an initial topic. The examiner takes charge of generating questions within a given topic. Then, the question will be sent to the pair of LLMs under test to produce the responses. Next, the judge compares the responses from the two LLMs and decides the winner for this question. Considering the responses for the current question, the eval controller will determine to deepen the current question if the current question results in a close tie or terminate the search along the current question if it is clear to decide the winner. We employ a breadth-first search strategy for this process. When we generate the next question, the eval controller also guarantees the diversity and reliability of the questions. Eventually, the nodes traversing over the tree will be aggregated to produce a comprehensive score for the evaluation.
TreeEval can be viewed as a tree-planning framework using LLMs as heuristic for hosting the evaluation session. By constructing a tree structure, we can differentiate the abilities of a pair of LLMs using as few as questions. Moreover, the test questions are automatically generated during each evaluation, which prevents the benchmark leakage. In the tree, the root node is initialized by the identified topic for the session. Each node contains a topic and a question within the topic. The connection between nodes indicates inheritance relationships between the questions, that is, the deepen question is derived from its parent question. The deeper the tree is, the more similar abilities the two LLMs under test hold. For sibling nodes in the tree, which are derived from the same parent node, they are designed to cover distinct topics but belong to the same superior topic.
3.2 TreeEval Modules
@@
Examiner. The examiner is a LLM-based module, which takes charge of generating exam questions that are able to cover diverse topics. Following Bai et al. (2023), we pre-define a set of topics as the scope of evaluation.
As the initialization of an evaluation session, we randomly sample a topic from the pre-defined topic set, which is denoted as $\mathcal{FC}_{\text{pre-define}}$. Given a topic, the examiner is requested to craft a question that related to it via a prompt with the consideration of the coherence to the topic and the required format of the question. The detailed instruction is displayed in Appendix 8.1.
@@
Judge. Previous studies Wang et al. (2023a) conduct pare-wise comparison and identify the superior responses among two evaluated LLMs, which has advantage in providing more nuanced assessment. Following these studies, we consult a pair of LLMs with the same question. The detailed instruction is displayed in Appendix 8.2. After the responses have been produced via the LLMs, another LLM performs as the judge to the responses.
@@
Inspired by the Tree-of-Thought Long (2023), where a controller produces the next thought step, we let Eval Controller arrange the follow-up evaluation according to $\mathcal{M}_{t}$. On the one hand, it prepares the follow-up topics $\mathcal{FC}_{t}$ based on $\{C_{t},A^{1}_{t},A^{2}_{t}\}\in\mathcal{M}_{t}$ for any of its child nodes in advance. On the other hand, it determines $Q_{t+1}$ based on the $\mathcal{FC}_{t}$ and $\{Q_{1},Q_{2},...,Q_{t}\}\in\mathcal{M}_{t}$ if the $t$-th node is the parent node at $t+1$ time step. We next describe the above two steps in detail:
  • Step One: Sample topics from the responses of the previous question: $\mathcal{FC}_{t}\sim\text{NER}(A_{t})$ [^2] The detailed instruction could be found in Appendix 8.3. This works better when the Named Entity Recognition (NER) tool is built upon a LLM as some relevant entities could be revised via the model instead of solely being extracted Wang et al. (2023b).
    We sample candidate topics from both $A_{t}^{1}$ and $A_{t}^{2}$ then merge them together, which results in a set of candidate topics $\tilde{\mathcal{FC}}_{t}$ as the follow-up topics of $t$-th node. However, this may produce some candidates that are repeated. To avoid this, we first measure the similarity between $C^{i}_{t}\in\tilde{\mathcal{FC}}_{t}$ and $C_{t}$ by computing the Cosine Similarity of their encoded vector representation Zhang et al. (2023a), which is denoted as $\text{Sim}(C^{i}_{t},C_{t})$. Then, we iteratively push out $C^{i}_{t}$ with the largest score. Next, we update the similarity scores of the rest topic $C^{j}_{t}$ by subtracting the similarity score of $C^{j}_{t}\in\tilde{\mathcal{FC}}_{t}\setminus C^{i}_{t}$ and $C^{i}_{t}$, which is to decrease the possibility of retrieving similar topics. This procedure continues until we have pushed out $k$ topics as $\mathcal{FC}_{t}$ for the follow-up question generation.
@@
  • After we have generated the sibling nodes of a parent node, we revisit the the scores of these siblings. If it shows a dominate score over all these siblings, this indicates that we have a winner for this branch. Hence we stop further search for any of these sibling nodes.
@@ 3.3 Score Aggregator
4 Experiments
LLMs MMLU ${}^{\star}$ BBH ${}^{\star}$ AlpacaEval ${}^{\dagger}$ MT-bench ${}^{\dagger}$ AlpacaEval2.0 ${}^{\dagger}$ TreeEval(Ours)
Acc Acc Win-Rate score Win-Rate #Q Score
Mistral-7B-Instruct-v0.2 $70.6$ $46.4$ $92.78$ $8.30$ $14.72$- $2.50$
Yi-34B-Chat $73.46$ $71.74$ $94.08$ $8.65$ $29.66$ $31.67$ $3.48$
xwinlm-13b-v0.1 $56.6$ $37.58$ $91.76$ $7.34$ $17.43$ $62.33$ $2.67$
Vicuna-33b-v1.3 $59.2$ $52.0$ $88.99$ $7.12$ $12.71$ $41.33$ $1.61$
WizardLM-13B-V1.2 $52.7$ $40.12$ $89.17$ $7.2$ $12.03$ $44.67$ $1.10$
zephyr-7b-beta $61.4$ $42.72$ $90.60$ $7.34$ $10.99$ $45.67$ $2.19$
@@ 4 Experiments
Table 1: Comparison of LLMs across various evaluation methods. $\star$” denotes we re-implement MMLU and BBH benchmarks Chia et al. (2023), calculating results in both $5$-shot and $3$-shot contexts. $\dagger$” denotes we directly take results from the respective leader-boards from MT-bench, AlpacaEval, and AlpacaEval2.0. “#Q” denotes the number of questions used for evaluation. We report the correlation of rankings obtained through different methods with those from AlpacaEval2.0, using $\tau$ for the Kendall correlation coefficient KENDALL (1938) and $\rho$ for the Spearman correlation coefficient Spearman (1904).
4.1 Experimental Setup
@@
Evaluated LLMs. We consider the following open-source LLMs as evaluated LLMs, encompassing two $7$B models, two $13$B models, and two $33$B models. They are either derived from LLaMA Touvron et al. (2023); et al. (2023a) or trained from scratch using the LLaMA architecture and some of them exhibit remarkably similar performances according to the open-source LLM leader-board [^3] https://tatsu-lab.github.io/alpaca_eval/.
In particular, Yi-34B-Chat 01.AI (2023) is a pioneering product from 01.AI. It is built on a large-scale multilingual dataset. Xwin-LM-13B-V0.1 Team (2023b) is built on the foundation of LLaMA2-13B, tuned through SFT and RLHF. Mistral-7B-Instruct-v0.2 Jiang et al. (2023) is tuned through SFT on the Mistral-7B model, which is built with the LLaMA architecture. Vicuna-33b-v1.3 Zheng et al. (2023a) originates from the LLaMA-33b model and is fine-tuned using user-shared dialogues from ShareGPT. WizardLM-13B-V1.2 Xu et al. (2023), based on the LLaMA2-13b model, is fine-tuned with enhanced instruction data using the Evol-Instruct. Zephyr-7B-beta Tunstall et al. (2023b), derived from the Mistral-7B model, is algined with SFT and DPO methods.
Comparable Evaluation Methods. We compare TreeEval with a series of existing evaluation methods. We include multiple methods with benchmark paradigm: MMLU Hendrycks et al. (2021), and Big-Bench Hard Suzgun et al. (2022) (BBH). Meanwhile, we include multiple methods treating LLMs as judges: AlpacaEval, AlpacaEval2.0 Li et al. (2023b), and MT-Bench Zheng et al. (2023a), which all choose ChatGPT as the judge but differ in their handling of dialogue scenarios. Specifically, MT-Bench is designed to assess multi-turn dialogues, whereas AlpacaEval focuses solely on evaluating single-turn interactions.
@@
Implementation Details. We utilize GPT-4-0613 as our examiner and deploy the model using FastChat Zheng et al. (2023a). We set the temperature parameter to $1$, which facilitates to generate variant questions. Furthermore, we set $T$ and $k$ as $3$. $\alpha$, $\beta$, and $\gamma$ are set as $1$, $1$, $0.4$, respectively. To mitigate the randomness of the experiments, we repeat $3$ times and calculate the average scores. We choose Mistral-7B-Instruct-v0.2 as the reference for pairwise comparison due to its moderate performance shown in public leaderboard.
4.2 Performance of TreeEval

### v2

@@ Abstract
Recently, numerous new benchmarks have been established to evaluate the performance of large language models (LLMs) via either computing a holistic score or employing another LLM as a judge. However, these approaches suffer from data leakage due to the open access of the benchmark and inflexible evaluation process. To address this issue, we introduce TreeEval, a benchmark-free evaluation method for LLMs that let a high-performance LLM host an irreproducible evaluation session and essentially avoids the data leakage. Moreover, this LLM performs as an examiner to raise up a series of questions under a topic with a tree planing strategy, which considers the current evaluation status to decide the next question generation and ensures the completeness and efficiency of the evaluation process. We evaluate $6$ models of different parameter sizes, including $7$B, $13$B, and $33$B, and ultimately achieved the highest correlation coefficient with AlpacaEval2.0 using only around $45$ questions. We also conduct more analysis to show the robustness and reliability of TreeEval. Our code can be accessed via the provided URL [^1] https://github.com/Ashura5/TreeEval Due to AAAI’s format limitations, the full version with appendices is available at (Li, Lan, and Yang 2024)..
Introduction
@@
The recent surge in Large Language Models (LLMs) has been significant, transitioning from closed-source (OpenAI 2023; Team 2023a) to open-source (Touvron et al. 2023; et al. 2023a; Jiang et al. 2023) models. Various Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) techniques have been proposed to further enhance the performance of LLMs (Taori et al. 2023; Chiang et al. 2023; Bai et al. 2022; Ouyang et al. 2022; Tunstall et al. 2023a). These LLMs demonstrate capabilities to address diverse tasks and are widely utilized in both academic and industrial fields. While human evaluation is intuitive for assessing the performance of LLMs, it is time-consuming and susceptible to unexpected bias (Zheng et al. 2023b; Wang et al. 2024). Thus, investigating automatic evaluation approaches for LLMs becomes crucial.
To date, numerous automatic evaluation methods have been proposed. One approach involves annotating benchmark datasets, such as MMLU and BBH (Hendrycks et al. 2021; Suzgun et al. 2022), to test various capabilities of an LLM. The performance is assessed by checking the overlap between annotated answers and generated answers, producing a holistic score to indicate the LLM’s performance. We refer to this category of evaluation methods as the benchmark paradigm. However, the holistic score can be inflexible for measuring the quality of LLM outputs since token mismatches do not necessarily indicate incorrect answers.
With the advent of high-performance LLMs, another approach leverages them to simulate human evaluation. This involves providing the evaluated LLM with predefined benchmark questions and using another LLM, such as GPT-4, to judge its responses (Zheng et al. 2023a; Li et al. 2023b; Bai et al. 2023; Wang et al. 2023a; Zhang et al. 2023b; Wang et al. 2023c; Li et al. 2023a; Zhu, Wang, and Wang 2023). We refer to this category of evaluation methods as the LLM-as-judge paradigm. However, this evaluation approach can also introduce additional biases, including positional bias (Wang et al. 2023a), verbosity bias (Saito et al. 2023), and style bias (Wu and Aji 2023). Positional bias refers to the tendency to assign higher scores to answers based on their specific positions. Verbosity bias indicates that large language models often prefer more verbose answers, even if these longer responses are not necessarily of higher quality than shorter ones. Style bias manifests in the inclination of large language models to favor answers that match their own generated style, such as giving lower scores to correct responses with spelling errors, since LLMs rarely produce content with spelling mistakes.
@@
Despite enabling automatic evaluation with standard pipelines, both the benchmark and LLM-as-judge paradigms face significant data leakage issues. The extensive training data used in LLM development, considered a valuable asset by many closed and even open-source models, can easily lead to benchmark data leakage, severely biasing evaluation results (Zhou et al. 2023b). To solve this issue, we propose a novel evaluation paradigm, which takes an LLM as an examiner to raise questions. The examiner should produce different evaluation session for each time which makes it hard to duplicate the evaluation questions and protect the evaluation benchmark from disclosure for fine-tuning and pre-training an LLM deliberately. However, simply adopting an LLM as examiner would lead to arbitrary evaluation question generation without a goal. Designing such a benchmark-free evaluation method need take the following aspects into consideration: (1) Similar as the question in a benchmark (Taori et al. 2023; Zheng et al. 2023a), the generated questions should be derived from certain topics, which ensures the scope of the evaluation. (2) Drawing inspiration from the interview, within a topic, the examiner should generate a line of questions that are diverse to cover different knowledge rather than producing a single question. (3) The generation procedure should be flexible enough to generate mutually connected questions and control the difficulty level of these questions. When the current line of question cannot distinguish two LLMs, more difficult questions should be raised up. Otherwise, the evaluation could be terminated immediately.
To this end, we propose TreeEval, which is a benchmark-free evaluation of the knowledge implication and question-answering capabilities of LLM through tree planning. The line of questions within a topic for evaluation are organized in a tree, where each node contains a question. In the process of constructing a tree, we repeatedly revisit the status of the current tree and generate the next node until the tree is enough to differentiate two LLMs. The difference between our evaluation method and previous paradigms can be found in Figure 1. To verify the effect of our method, we evaluate multiple LLMs. The results demonstrate that our method shows similar ranking as AlpacaEval2.0 in LLM-as-judge paradigm with only $45$ questions in average for each round of evaluation. Further analysis shows our advantages in measuring fine-grained capabilities and conducting robust comparison for LLMs.
@@
  • We compare with a set of automatic evaluation baselines, and find that our TreeEval achieves the highest correlation coefficient with AlpacaEval2.0.
Related Work
Methods of LLM Evaluation
Due to the explosive growth and rapid update of LLMs, a significant challenge is to conduct accurate and comprehensive evaluation for them (Chang et al. 2023). Early studies leverage open-ended question answering datasets and math word problems as the evaluation benchmarks (Touvron et al. 2023; et al. 2023b; Chen et al. 2024) to evaluate the commonsense knowledge and reasoning capabilities of LLMs. Subsequently, more benchmark datasets like MMLU (Hendrycks et al. 2021), AGIEval (Zhong et al. 2023), IFEval (Zhou et al. 2023a) have been elaborately designed to gauge diverse abilities of LLMs. Some studies (Wang et al. 2023a, a; Saha et al. 2023) go beyond standard evaluation metrics. They evaluate the quality and accuracy of predicted results through human annotation, which is able to provide a more comprehensive feedback. With the emergence of high-performance LLMs like GPT-4 (OpenAI 2023), Gemini Pro (Team 2023a), more recent studies start to utilize them to simulate the human evaluation process. In this realm, PandaLM (Wang et al. 2023c) strives to provide reproducible and automated comparison between various LLMs by training a LLM as the judge. GPTScore (Fu et al. 2023) and G-Eval (Liu et al. 2023) utilize GPT-3 and GPT-4 as the judge to evaluate the LLMs with incorporation of in-context learning and chain-of-thought strategies. The above methods rely heavily on a well-organized benchmark dataset. However, there have been some recent works focusing on data leakage of LLM reviews. (Zhu et al. 2024) proposed a method based on DAG to dynamically generate samples to evaluate LLM reasoning capabilities during the evaluation process. And our method is benchmark-free and has LLMs performing as the examiner to evaluate other models’ knowledge entailment and question answering capabilities.
Data Leakage of LLM Evaluation
@@
Methodology
Figure 2: TreeEval system with an illustrative tree for evaluation. The left section contains the components and their workflow in TreeEval. The right section displays a constructed tree within topic Technology and Communication for evaluation (the leaf nodes are shown in red boxes), where each node denotes a question annotated with its topic and evaluation score. We further display the generated questions of the tree in the (Appendix Eval Controller Example).
Overall Architecture
Figure 2 shows the overall structure of TreeEval. TreeEval organizes evaluations in a tree format, using components such as an Examiner, a Judge, and an Eval Controller. After the tree is built, an Aggregator compiles the scores. This framework allows for benchmark-free evaluation of LLMs through tree planning. Here’s how it works:
  1. 1.
    Session Setup: For each evaluation session, we choose two LLMs and start with an initial topic.
  2. 2.
    Question Generation: The Examiner generates questions within this topic.
  3. 3.
    Response Collection: These questions are sent to the LLMs, and their responses are collected.
  4. 4.
    Response Evaluation: The Judge compares the responses and decides the winner for each question.
  5. 5.
    Evaluation Control: If the responses are closely matched, the Eval Controller deepens the question. If a clear winner is found, the process moves to a new question. This follows a breadth-first search strategy, ensuring diverse and reliable questions.
  6. 6.
    Score Aggregation: Finally, the Aggregator compiles the scores from all nodes in the tree to produce a comprehensive evaluation score.
TreeEval uses a tree structure to evaluate LLMs, minimizing the number of questions needed. Questions are generated automatically, preventing benchmark leakage. The root node starts with the session’s topic, and each node represents a question within that topic. Connections between nodes show how questions evolve. Deeper nodes indicate more similar abilities between the LLMs. Sibling nodes, derived from the same parent, cover different subtopics of the same main topic.
TreeEval Modules
@@
Examiner. The examiner is a LLM-based module, which takes charge of generating exam questions that are able to cover diverse topics. Following (Bai et al. 2023), we pre-define a set of topics as the scope of evaluation.
As the initialization of an evaluation session, we randomly sample a topic from the pre-defined topic set, which is denoted as $\mathcal{FC}_{\text{pre-define}}$. Given a topic, the examiner is requested to craft a question that related to it via a prompt with the consideration of the coherence to the topic and the required format of the question. The detailed instruction is displayed in (Appendix Prompt for Examiner).
@@
Judge. Previous studies (Wang et al. 2023a) conduct pairwise comparison and identify the superior responses among two evaluated LLMs, which has advantage in providing more nuanced assessment. Following these studies, we consult a pair of LLMs with the same question. The detailed instruction is displayed in (Appendix Prompt for Judge). After the responses have been produced via the LLMs, another LLM performs as the judge to the responses.
@@
Inspired by the Tree-of-Thought (Long 2023), where a controller produces the next thought step, we let Eval Controller arrange the follow-up evaluation according to $\mathcal{M}_{t}$. On the one hand, it prepares the follow-up topics $\mathcal{FC}_{t}$ based on $\{C_{t},A^{1}_{t},A^{2}_{t}\}\in\mathcal{M}_{t}$ for any of its child nodes in advance. On the other hand, it determines $Q_{t+1}$ based on the $\mathcal{FC}_{t}$ and $\{Q_{1},Q_{2},...,Q_{t}\}\in\mathcal{M}_{t}$ if the $t$-th node is the parent node at $t+1$ time step. We next describe the above two steps in detail:
  • Step One: Sample topics from the responses of the previous question: $\mathcal{FC}_{t}\sim\text{NER}(A_{t})$ [^2] The detailed instruction could be found in (Appendix Prompt for NER). This works better when the Named Entity Recognition (NER) tool is built upon a LLM as some relevant entities could be revised via the model instead of solely being extracted (Wang et al. 2023b).
    We sample candidate topics from both $A_{t}^{1}$ and $A_{t}^{2}$ then merge them together, which results in a set of candidate topics $\tilde{\mathcal{FC}}_{t}$ as the follow-up topics of $t$-th node. However, this may produce some candidates that are repeated. To avoid this, we first measure the similarity between $C^{i}_{t}\in\tilde{\mathcal{FC}}_{t}$ and $C_{t}$ by computing the Cosine Similarity of their encoded vector representation (Zhang et al. 2023a), which is denoted as $\text{Sim}(C^{i}_{t},C_{t})$. Then, we iteratively push out $C^{i}_{t}$ with the largest score. Next, we update the similarity scores of the rest topic $C^{j}_{t}$ by subtracting the similarity score of $C^{j}_{t}\in\tilde{\mathcal{FC}}_{t}\setminus C^{i}_{t}$ and $C^{i}_{t}$, which is to decrease the possibility of retrieving similar topics. This procedure continues until we have pushed out $k$ topics as $\mathcal{FC}_{t}$ for the follow-up question generation.
@@
  • After we have generated the sibling nodes of a parent node, we revisit the scores of these siblings. If it shows a dominate score over all these siblings, this indicates that we have a winner for this branch. Hence we stop further search for any of these sibling nodes.
@@
Experiments
LLMs MMLU BBH AlpacaEval MT-bench AlpacaEval2.0 TreeEval(Ours)
Acc Acc Win-Rate score Win-Rate #Q Score(var)
Mistral-7B-Instruct-v0.2 $70.6$ $46.4$ $92.78$ $8.30$ $14.72$- $2.50(0.000)$
Yi-34B-Chat $73.46$ $71.74$ $94.08$ $8.65$ $27.19$ $31.67$ $3.48(0.011)$
xwinlm-13b-v0.1 $56.6$ $37.58$ $91.76$ $7.34$ $17.43$ $62.33$ $2.67(0.000)$
WizardLM-13B-V1.2 $52.7$ $40.12$ $89.17$ $7.2$ $12.03$ $44.67$ $1.10(0.070)$
zephyr-7b-beta $61.4$ $42.72$ $90.60$ $7.34$ $10.99$ $45.67$ $2.19(0.003)$
Vicuna-33b-v1.3 $59.2$ $52.0$ $88.99$ $7.12$ $12.71$ $41.33$ $1.61(0.044)$
@@
Table 1: Comparison of LLMs across various evaluation methods. $\star$” denotes we re-implement MMLU and BBH benchmarks (Chia et al. 2023), calculating results in both $5$-shot and $3$-shot contexts. $\dagger$” denotes we directly take results from the respective leader-boards from MT-bench, AlpacaEval, and AlpacaEval2.0. “#Q” denotes the number of questions used for evaluation. We report the correlation of rankings obtained through different methods with those from AlpacaEval2.0, using $\tau$ for the Kendall correlation coefficient (KENDALL 1938) and $\rho$ for the Spearman correlation coefficient (Spearman 1904).
Experimental Setup
@@
Evaluated LLMs. We evaluated the following open-source LLMs, including two 7B models, two 13B models, and two 33B models. These models are either derived from LLaMA (Touvron et al. 2023; et al. 2023a) or trained from scratch using the LLaMA architecture, and some show similar performance according to the open-source LLM leaderboard [^3] https://tatsu-lab.github.io/alpaca˙eval/.
  • Yi-34B-Chat (01.AI 2023) is a product from 01.AI, built on a large-scale multilingual dataset.
  • Xwin-LM-13B-V0.1 (Team 2023b) is based on LLaMA2-13B and tuned through SFT and RLHF.
  • Mistral-7B-Instruct-v0.2 (Jiang et al. 2023) is tuned on the Mistral-7B model, built with the LLaMA architecture.
  • Vicuna-33B-v1.3 (Zheng et al. 2023a) originates from LLaMA-33B and is fine-tuned using dialogues from ShareGPT.
  • WizardLM-13B-V1.2 (Xu et al. 2023) is based on LLaMA2-13B and fine-tuned with enhanced instruction data using Evol-Instruct.
  • Zephyr-7B-beta (Tunstall et al. 2023b) is derived from Mistral-7B and aligned using SFT and DPO methods.
Comparable Evaluation Methods. We compare TreeEval with several existing methods, including:
  • Benchmark Paradigm:
    • MMLU (Hendrycks et al. 2021)
    • Big-Bench Hard (BBH) (Suzgun et al. 2022)
  • LLMs as Judges:
    • AlpacaEval and AlpacaEval2.0 (Li et al. 2023b)
    • MT-Bench (Zheng et al. 2023a)
AlpacaEval and AlpacaEval2.0 use ChatGPT as the judge for single-turn interactions, while MT-Bench focuses on multi-turn dialogues.
@@
Implementation Details. We use GPT-4-0613 as the examiner, deployed with FastChat (Zheng et al. 2023a). The temperature is set to 1 to generate varied questions. We set $T$ and $k$ to 3. Parameters $\alpha$, $\beta$, and $\gamma$ are set to 1, 1, and 0.4, respectively. To reduce randomness, we repeat the experiments three times, average the scores, and additionally report the variance across runs to demonstrate that our method enables more stable comparisons between models with similar performance. We use Mistral-7B-Instruct-v0.2 as the reference model for pairwise comparison due to its moderate performance on public leaderboards.
Performance of TreeEval